# AI-Generated Images Introduce Invisible Relevance Bias to Text-Image Retrieval

### v3

@@ Abstract.
With the advancement of generation models, AI-generated content (AIGC) is becoming more realistic and less discernible, flooding the Internet. A recent study suggests that this phenomenon causes source bias in text retrieval for web search. Specifically, neural retrieval models tend to rank generated texts higher than human-written texts.
In this paper, we extend the study of this bias to cross-modal retrieval. Firstly, we successfully construct a suitable benchmark to explore the existence of the bias. In this benchmark, AI-generated images possess visual semantics sufficiently similar to real images. Subsequent extensive experiments on this benchmark reveal that AI-generated images introduce an invisible relevance bias to text-image retrieval models. Specifically, our experiments show that text-image retrieval models tend to rank the AI-generated images higher than the real images, even though the AI-generated images do not exhibit more visually relevant features to the query than real images. This invisible relevance bias is prevalent across retrieval models with varying training data and architectures, encompassing models trained from scratch to those pre-trained on extensive image-text pairs, and spanning dual-encoder retrieval models to fusion-encoder retrieval models. Furthermore, our subsequent exploration reveals that the inclusion of AI-generated images in the training data of the retrieval models exacerbates the invisible relevance bias. The above phenomenon triggers a vicious cycle in which AI-generated images have a higher chance of being obtained from massive amounts of data, which makes them more likely to be mixed into the training of retrieval models and such training makes the invisible relevance bias become more and more serious.
To elucidate the potential causes of invisible relevance and address the aforementioned issues, we introduce an effective training method aimed at alleviating the invisible relevance bias. Subsequently, we apply our proposed debiasing method to retroactively identify the causes of invisible relevance, revealing that the AI-generated images induce the image encoder to embed additional information into their representation. This information exhibits a certain consistency across generated images with different semantics and can make the retriever estimate a higher relevance score. We conduct experiments to substantiate this assertion.
@@ 1. Introduction
With the advancement of generation models, AI-generated content (AIGC) has become increasingly realistic and less distinguishable (Yang et al., 2023; Brown et al., 2020). The utilization of AI for content generation has significantly transformed the landscape of content creation. It has not only reduced the cost of content generation but has also enhanced efficiency, leading to a rapid influx of a substantial volume of AI-generated content onto the Internet (Ai et al., 2023; Dai et al., 2023).
Information retrieval (IR) is an important way for people to obtain the target information from massive Internet information (Manning, 2009). However, the rapid proliferation of AI-generated content (AIGC) presents a significant new challenge to IR: as the Internet becomes increasingly inundated with AI-generated contents, how will these contents influence the ranking results of search? In response to this challenge, (Dai et al., 2023) conducted a study on text modality and found that neural information retrieval models exhibit a preference for texts generated by Large Language Models. They refer to this category of biases in neural retrieval models towards the LLM-generated text as the source bias. However, beyond textual information, the Internet is replete with a substantial number of images, serving as crucial sources for IR systems. Building on this, our paper extends the investigation of source bias in AI-generated content to text-image retrieval models. A pivotal question emerges: What impact will AI-generated images have on existing text-image retrieval models?
@@
Another important point that is not considered in previous work is that AI-generated content does not only have an impact on the inference stage of the retrieval models. Due to their wide distribution on the Internet, they are very likely to be mixed into the training data of retrieval models in the future. Therefore, we further mix the AI-generated images into the training data of the retrieval model and explore the rules of its impact on the retrieval models by adjusting the mixing ratio. Our experiments show that as the mixing ratio of AI-generated images in training data increases, the invisible relevance bias becomes more serious (§ 3.4), the retrieval model exhibits a greater inclination to rank generated images higher. This phenomenon reveals that invisible relevance bias causes the generated images to have a higher chance of being obtained from massive data, which makes them more likely to be mixed into the training of generation and retrieval models and causes the model collapse (Shumailov et al., 2023), leading to more serious bias and forming a vicious cycle. (shown in Figure 1)
@@ 2. Benchmark Construction
This paper focuses on revealing the potential impact of AI-generated images on the ranking results of text-image retrieval systems. The prerequisite for all this is to construct a retrieval database consisting of real images and AI-generated images. This database should meet four critical requirements. (-$\mathcal{R}\mbox{-}1$) Firstly, from the perspective of fair assessment, a reasonable scenario for assessing the potential bias requires that the generated images and the real images must have sufficiently similar visual semantics. This can avoid increasing or decreasing some semantic associations between the generated images and the query caused by image generation. That is, the IR model preferring (or rejecting) an AI-generated image that is more (or less) semantically relevant to the query than the real image cannot prove the existence (or nonexistence) of the bias. (-$\mathcal{R}\mbox{-}2$) Secondly, retrieval performance on the database containing only generated images should not change significantly compared to retrieval performance on real images only. This can further ensure that no additional visual semantics relevant (or irrelevant) to the query are introduced during the image generation. (-$\mathcal{R}\mbox{-}3$) Thirdly, the image generation pipeline should be in line with the most common scenario of content generation that receives a description as prompt and generates the texts or images, so that the obtained bias assessment results can be more consistent with the actual scene. (-$\mathcal{R}\mbox{-}4$) Fourthly, the number of generated images and real images should be equal to achieve fair comparison.
Table 1. Compare generation methods based on the four requirements outlined in Section 2. Similarity is the cosine between embeddings of real and generated images encoded by OpenClip. Retrieval Performance is the difference in retrieval performance of BEIT-3 on the databases only containing generated images and only containing real images respectively.
@@ 2.1. Image Generation
Generation Based on Merged Caption. Since a single caption may not contain the complete visual semantics of the image, we propose to merge five captions to obtain a relatively complete semantic description of the image. Specifically, for each real image ($I_{r}$), we use gpt-3.5-turbo to combine the five captions to get a newly merged caption. And input the merged caption to a stable diffusion model $\mathbf{M}$ to generate the image ($I_{g}$), which can be described as:
$C_{m}=\textrm{GPT-3.5}(C_{1},C_{2},...,C_{n}),\quad I^{g}=\mathbf{M}(C_{m}).$
The intention of this design is that different captions of an image can be the descriptions from different perspectives of the image. Therefore, merging these captions can obtain an overall description of the image, thereby making stable diffusion generate images that have sufficiently similar visual semantics to the real image. Besides, this also avoids the relevance bias in assessment caused by the generated image being overly inclined to one certain caption.
@@
Generation Based on Image Rewriting. In addition to the merged caption $C_{m}$, the real image $I_{r}$ is also used as the initial image to condition the generation of the new image. Generation constrained by the input real image can output the generated image that has higher similarity to the real image. But it also puts forward higher requirements for the multi-modal understanding ability of the model.
@@
Generation Based on Image Over Sampling and Selection. To further narrow the semantic similarity between the generated images and real images, we propose an image over-sampling strategy. Specifically, for a real image $I^{r}$, we use stable diffusion to perform multiple times generation with different random seeds and get $n$ generated images $\{{I^{g}_{1}},{I^{g}_{2}},...,{I^{g}_{n}}\}$. Then, we use the vision encoder $v(\cdot;\theta)$ of a powerful open source pre-trained vision-language model [^2] https://huggingface.co/laion/CLIP-ViT-H-14-laion2B-s32B-b79K to get the embedding $\bm{e_{r}}$ for $I^{r}$ and set of embeddings $\bm{E}=\{\bm{{e^{g}_{1}}},\bm{{e^{g}_{2}}},...,\bm{{e^{g}_{n}}}\}$ for $\{{I^{g}_{1}},{I^{g}_{2}},...,{I^{g}_{n}}\}$. We calculate the cosine similarity between each embedding in $\bm{E}$ and $\bm{e_{r}}$ to get the similarity between the generated images and the real image $I^{r}$. Finally, we select the generated image with the maximum similarity between $I^{r}$ as the final generated image ${I}^{g}$ for the real image:
(1) $I^{g}=\arg\max_{i}\{cosine(\bm{e^{r}},\bm{e^{g}_{i}})|\bm{{e^{g}_{i}}}\in\bm{E}\}.$
@@
Comparison between Different Generation Methods. The comparison between different generation methods under the above-mentioned four requirements is shown in Table 1. Taking these results into account, we choose generation based on merged caption and image over-sampling and selection as the image generation method to construct our benchmark. Using this method, we can get an AI-generated image for each real image in Flicker30k and MSCOCO. For the convenience of description, in the following contents, Flicker30k and MSCOCO indicate the datasets that only contain real images, Flicker30k (AI) and MSCOCO (AI) indicate the datasets that only contain AI-generated images, Flicker30k+AI and MSCOCO+AI indicate the datasets that consist of both real and AI-generated images..
2.2. Assessing the Quality of Generated Images
@@
Ensuring sufficiently similar visual semantics between generated and real images to avoid increasing or decreasing relevance to the query is a prerequisite for assessing the bias. In this section, we propose two methods to evaluate this. Firstly, we perform retrieval on the databases that only contain real images and only contain generated images and count the retrieval performance respectively. The intention for this is that if the generated images have more semantics relevant to their corresponding queries than the real images, distinguishing images becomes easier, resulting in significantly higher retrieval performance on the generated images database than on the real images database. Secondly, we introduce human evaluation to further evaluate the semantic information from human vision. Specifically, we let humans determine whether the generated images have more visual semantics relevant to the queries compared to the real images.
@@ 2.2.1. Retrieval Performance
Table 2. Retrieval performance (NDCG) on the databases that only contain real images and only contain AI-generated images. Performance on AI-generated images is not significantly better than the performance on real images can demonstrate the success of our benchmark construction.
@@
The experimental results in Table 2 show that retrieval performance on the database containing only AI-generated images is not significantly greater than retrieval performance on the database containing only real images. It indicates that the AI-generated images in our benchmark do not introduce more visual semantics relevant to the query. Specifically, we use three open-source and powerful vision-language pre-trained models including FLAVA (Singh et al., 2022), ALIGN (Jia et al., 2021), and BEIT-3 (Wang et al., 2022) to perform retrieval on the databases that only contain real images and only contain generated images respectively. Since text-image retrieval datasets are composed of real images, the training on these images will introduce additional bias. Considering the excellent zero-shot text-image retrieval performance of the three models (Wang et al., 2022), we use these models directly for retrieval in zero-shot setting.
@@ 2.2.2. Human Evaluation
Table 3 shows humans think that in most samples of our benchmark, real images have more or equal visual semantics relevant to the queries than AI-generated images. This further guarantees the fairness of our assessment of invisible relevance bias. If the AI-generated images do not have more relevant visual semantics than real images, while the text-image retrieval model still tends to rank AI-generated images higher than the real images, the invisible relevance bias does exist. Specifically, we invite five humans with master’s degrees to participate in the evaluation. The triple consisting of a caption (i.e., the query), a real image, and its corresponding AI-generated image is the basic unit in this evaluation. We ask the five humans to select which image (real or AI-generated) has more relevant semantics to the caption. We count the proportion of selections made by humans on our benchmark and the results are shown in Table 3.
@@ 3. Bias Assessment
Based on the benchmark constructed above, we can assess the impact of AI-generated images on text-image retrieval. We not only study the impact when AI-generated images appear in the database but also further discuss the problems caused when AI-generated images are mixed into the training data of retrieval models.
@@ 3.2. Experimental Settings and Metrics
As the neural networks tend to fit the data domian in training (Xu et al., 2022), our assessment is performed under the out-of-domain setting to try to mitigate potential bias introduced by the domain of the training data. Specifically, for the models that need to train from scratch on supervised text-image pairs (NAAF and VSE), we train them on Flicker30k (MSCOCO) and evaluate their performance on MSCOCO (Flicker30k). For the models that have been pre-trained on massive real text-image pairs and show excellent zero-shot performance in text-image retrieval, we directly use these pre-trained models to perform retrieval on the test datasets. An exception is that even though VILT has been pre-trained, it needs to be combined with a specific multi-layer perceptron to complete the text-image matching task in text-image retrieval. So we fine-tune VILT on supervised text-image retrieval datasets just like NAAF and VSE.
@@ 3.3. Invisible Relevance Bias
The experimental results are shown in Table 4. Overall, invisible relevance bias widely exists in text-image retrieval models, that is, text-image retrieval models tend to rank AI-generated images higher than real images even though they have very similar semantics. Specifically, the following conclusions can be made: (1) The invisible relevance bias exists in both the models trained from scratch and the vision-language models that have been pre-trained on massive supervised text-image pairs. (2) The invisible relevance bias exists in both dual-encoder-based and fusion-encoder-based retrieval models. (3) The invisible relevance bias has a relatively greater impact on the Top-1 retrieved image. In the retrieved list, the Top-1 item is most likely to be clicked by users, which means that invisible relevance bias introduced by AI-generated images will have a huge impact on users’ actual search and click results.

### v4

# Invisible Relevance Bias: Text-Image Retrieval Models Prefer AI-Generated Images

@@ Abstract.
With the application of generation models, internet is increasingly inundated with AI-generated content (AIGC), causing both real and AI-generated content indexed in corpus for search. This paper explores the impact of AI-generated images on text-image search in this scenario. Firstly, we construct a benchmark consisting of both real and AI-generated images for this study. In this benchmark, AI-generated images possess visual semantics sufficiently similar to real images. Experiments on this benchmark reveal that text-image retrieval models tend to rank the AI-generated images higher than the real images, even though the AI-generated images do not exhibit more visually relevant semantics to the queries than real images. We call this bias as invisible relevance bias. This bias is detected across retrieval models with different training data and architectures. Further exploration reveals that mixing AI-generated images into the training data of retrieval models exacerbates the invisible relevance bias. These problems cause a vicious cycle in which AI-generated images have a higher chance of exposing from massive data, which makes them more likely to be mixed into the training of retrieval models and such training makes the invisible relevance bias more and more serious. To mitigate this bias and elucidate the potential causes of the bias, firstly, we propose an effective method to alleviate this bias. Subsequently, we apply our proposed debiasing method to retroactively identify the causes of this bias, revealing that the AI-generated images induce the image encoder to embed additional information into their representation. This information makes the retriever estimate a higher relevance score. We conduct experiments to support this assertion.
Findings in this paper reveal the potential impact of AI-generated images on retrieval and have implications for further research. Code is released at https://github.com/xsc1234/Invisible-Relevance-Bias.
Text-Image Retrieval, AIGC, Bias and Fairness
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USAbooktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington, DC, USAdoi: 10.1145/3626772.3657750isbn: 979-8-4007-0431-4/24/07ccs: Information systems Retrieval models and ranking
@@ 1. Introduction
With the advancement of generation models, the quality of AI-generated content (AIGC) has been increasingly improved (Yang et al., 2023; Brown et al., 2020). The utilization of AI for content generation has transformed the way of content creation. It not only reduces the cost of content generation but also enhances efficiency, leading to a rapid influx of large amounts of AI-generated content onto the internet (Ai et al., 2023; Dai et al., 2023).
Information retrieval (IR) is an important way for people to obtain the target information from massive data (Manning, 2009). However, the rapid proliferation of AI-generated content (AIGC) presents a significant new challenge to IR: As the internet becomes increasingly inundated with AI-generated content, the corpus for search contains both real and AI-generated content, so, how will AI-generated content influence the ranking of search? In response to this challenge, (Dai et al., 2023) conducted a study on text modality and found that neural information retrieval models exhibit a preference for texts generated by Large Language Models. They refer to this category of biases in neural retrieval models towards the LLM-generated text as the source bias. However, beyond textual information, the internet is replete with a substantial number of images, serving as crucial sources for IR systems. Based on this, our paper extends the investigation of source bias in AI-generated content to text-image retrieval models. A pivotal question emerges: What impact will AI-generated images have on existing text-image retrieval models?
@@
Another important point that is not considered in previous work is that AI-generated content does not only have an impact on the inference stage of the retrieval models. Due to the wide distribution of AI-generated images on the internet, they are very likely to be mixed into the training data of retrieval models in the future. Therefore, we further mix the AI-generated images into the training data of the retrieval model and explore the rules of its impact on the retrieval models by adjusting the mixing ratio. Our experiments show that as the mixing ratio of AI-generated images in training data increases, the invisible relevance bias becomes more serious (§ 3.4), and the retrieval model exhibits a greater inclination to rank generated images higher. This phenomenon reveals a vicious cycle (Figure 1): Invisible relevance bias causes the generated images to have a higher chance of being obtained from massive data, which makes them more likely to be mixed into the training of generation and retrieval models and causes the model collapse (Shumailov et al., 2023), leading to more serious bias and forming a cycle.
@@ 2. Benchmark Construction
This paper focuses on revealing the potential impact of AI-generated images on the ranking results of text-image retrieval systems. The prerequisite for all this is to construct a retrieval corpus consisting of real images and AI-generated images. This corpus should meet four critical requirements. (-$\mathcal{R}\mbox{-}1$) Firstly, from the perspective of fair assessment, a reasonable scenario for assessing the potential bias requires that the generated images and the real images must have sufficiently similar visual semantics. This can avoid increasing or decreasing some semantic associations between the generated images and the queries caused by image generation. That is, the IR model preferring (or rejecting) an AI-generated image that is more (or less) semantically relevant to the query than the real image cannot prove the existence (or nonexistence) of the bias. (-$\mathcal{R}\mbox{-}2$) Secondly, retrieval performance on the corpus containing only generated images should not change significantly compared to retrieval performance on real images only. This can further ensure that no additional visual semantics relevant (or irrelevant) to the query are introduced during the image generation. (-$\mathcal{R}\mbox{-}3$) Thirdly, the image generation pipeline should be in line with the most common scenario of content generation that receives a description as prompt and generates the texts or images, so that the obtained bias assessment results can be more consistent with the actual scene. (-$\mathcal{R}\mbox{-}4$) Fourthly, the number of generated images and real images should be equal to achieve fair comparison.
Table 1. Compare generation methods based on the four requirements outlined in Section 2. Similarity is the cosine between embeddings of real and generated images encoded by OpenClip. Retrieval Performance is the difference in retrieval performance of BEIT-3 on the corpora only containing generated images and only containing real images respectively.
@@ 2.1. Image Generation
Generation Based on Merged Caption. Since a single caption may not contain the complete visual semantics of the image, we propose to merge five captions to obtain a relatively complete semantic description of the image. Specifically, for each real image ($I_{r}$), we use gpt-3.5-turbo to combine the five captions to get a newly merged caption $C_{m}$. And input the merged caption to a stable diffusion model $\mathbf{M}$ to generate the image ($I_{g}$). The intention of this design is that different captions of an image can be the descriptions from different perspectives of the image. Therefore, merging these captions can obtain an overall description of the image, thereby making stable diffusion generate images that have sufficiently similar visual semantics to the real image.
@@
Generation Based on Image Rewriting. In addition to the merged caption $C_{m}$, the real image $I_{r}$ is also used as the initial image to condition the generation of the new image. Generation constrained by the input real image can output the generated image that has higher similarity to the real image. But it also requires the generation model to have higher multi-modal understanding ability.
@@
Generation Based on Image Over Sampling and Selection. To further narrow the semantic similarity between the generated images and real images, we propose an image over-sampling strategy. Specifically, for a real image $I^{r}$ we use stable diffusion to perform multiple times generation with different random seeds and get $n$ generated images $\{{I^{g}_{1}},{I^{g}_{2}},...,{I^{g}_{n}}\}$. Then, we use the vision encoder $v(\cdot;\theta)$ of a powerful open source pre-trained vision-language model [^2] https://huggingface.co/laion/CLIP-ViT-H-14-laion2B-s32B-b79K to get the embedding $\bm{e_{r}}$ for $I^{r}$ and set of embeddings $\bm{E}=\{\bm{{e^{g}_{1}}},\bm{{e^{g}_{2}}},...,\bm{{e^{g}_{n}}}\}$ for $\{{I^{g}_{1}},{I^{g}_{2}},...,{I^{g}_{n}}\}$. We calculate the cosine similarity between each embedding in $\bm{E}$ and $\bm{e_{r}}$ to get the similarity between the generated images and the real image $I^{r}$. Finally, we select the generated image with the maximum similarity between $I^{r}$ as the final generated image ${I}^{g}$ for the real image.
@@
Comparison between Different Generation Methods. The comparison between different generation methods under the above-mentioned four requirements is shown in Table 1. Taking these results into account, we choose generation based on merged caption and image over-sampling and selection as the image generation method to construct our benchmark. Using this method, we can get an AI-generated image for each real image in Flicker30k and MSCOCO. In the following content, Flicker30k and MSCOCO indicate the datasets that only contain real images, Flicker30k (AI) and MSCOCO (AI) indicate the datasets that only contain AI-generated images, Flicker30k+AI and MSCOCO+AI indicate the datasets that consist of both real and AI-generated images.
2.2. Assessing the Quality of Benchmark
@@
Ensuring sufficiently similar visual semantics between generated and real images to avoid increasing or decreasing relevance to the query is a prerequisite for assessing the bias. In this section, we propose two methods to evaluate this. Firstly, we perform retrieval on the corpora that only contain real images and only contain generated images and count the retrieval performance respectively. The intention for this is that if the generated images have more semantics relevant to their corresponding queries than the real images, distinguishing images becomes easier, resulting in significantly higher retrieval performance on the generated images corpus than on the real images corpus. Secondly, we introduce human evaluation to determine whether the generated images have more visual semantics relevant to the queries.
@@ 2.2.1. Retrieval Performance
Table 2. Retrieval performance (NDCG) on the corpora that only contain real images and only contain AI-generated images. Performance on AI-generated images is not significantly better than the performance on real images can demonstrate the success of our benchmark construction.
@@
The experimental results in Table 2 show that retrieval performance on the corpus containing only AI-generated images is not significantly greater than retrieval performance on the corpus containing only real images. It indicates that the AI-generated images in our benchmark do not introduce more visual semantics relevant to the queries. Specifically, we use three open-source and powerful vision-language pre-trained models including FLAVA (Singh et al., 2022), ALIGN (Jia et al., 2021), and BEIT-3 (Wang et al., 2022) to perform retrieval on the corpora that only contain real images and only contain generated images respectively. Since text-image retrieval datasets are composed of real images, the training on these images will introduce additional bias. Therefore, considering the excellent zero-shot text-image retrieval performance of the three models (Wang et al., 2022), we use these models directly for retrieval in zero-shot setting.
@@ 2.2.2. Human Evaluation
Table 3 shows humans think that in most samples of our benchmark, real images have more or equal visual semantics relevant to the queries than AI-generated images. This further guarantees the fairness of our assessment of invisible relevance bias. If the AI-generated images do not have more relevant visual semantics than real images, while the text-image retrieval model still tends to rank AI-generated images higher than the real images, the invisible relevance bias does exist. Specifically, we invite five humans with master’s degrees to participate in the evaluation. We randomly shuffle the dataset into five parts and assign each part randomly to the volunteers. Volunteers are shown with a triple consisting of a caption (i.e., the query), a real image, and its corresponding AI-generated image. We ask them to select which image (real or AI-generated) has more relevant semantics to the caption. We instruct volunteers how to visually judge the relevance between images and captions, and calculate the pass rate through cross-validation between two different volunteers. The formal human evaluation begins when each volunteer’s pass rate reaches 95%. We count the proportion of selections made by humans on our benchmark and the results are shown in Table 3.
@@ 3. Bias Assessment
Based on the benchmark constructed above, we can assess the impact of AI-generated images on text-image retrieval. We not only study the impact when AI-generated images appear in the corpus, but also further discuss the impact when AI-generated images are mixed into the training data of retrieval models.
@@ 3.2. Experimental Settings and Metrics
As the neural networks tend to fit the data domain in training (Xu et al., 2022), our assessment is performed under the out-of-domain setting to try to mitigate potential bias introduced by the domain of the training data. Specifically, for the models that need to train from scratch on supervised text-image pairs (NAAF and VSE), we train them on Flicker30k (MSCOCO) and evaluate their performance on MSCOCO (Flicker30k). For the models that have been pre-trained on massive real text-image pairs and show excellent zero-shot performance in text-image retrieval, we directly use these pre-trained models to perform retrieval on the test datasets. An exception is that even though VILT has been pre-trained, it needs to be combined with a specific multi-layer perceptron to complete the text-image matching task in text-image retrieval. So we fine-tune VILT on supervised text-image retrieval datasets just like NAAF and VSE.
@@ 3.3. Invisible Relevance Bias
The experimental results are shown in Table 4. Overall, invisible relevance bias widely exists in text-image retrieval models, that is, text-image retrieval models tend to rank AI-generated images higher than real images even though they have very similar visual semantics. Specifically, the following conclusions can be made: (1) The invisible relevance bias exists in both the models trained from scratch and the vision-language models that have been pre-trained on massive supervised text-image pairs. (2) The invisible relevance bias exists in both dual-encoder-based and fusion-encoder-based retrieval models. (3) The invisible relevance bias has a relatively greater impact on the Top-1 retrieved image. In the retrieved list, the Top-1 item is most likely to be clicked by users, which means that invisible relevance bias introduced by AI-generated images will have a huge impact on users’ actual search and click results.